Water is added to 25ml of 0.866M KNO3 solution until the volume of the solution is exactly 500 ml. What is the concentration of the final solution
1
Expert's answer
2020-03-02T06:55:04-0500

C=C0V0/V=0.866*25*10-3/500*10-3=0.0433M